### v1

# What Color Is It? A Text-Interference Multimodal Hallucination Benchmark

Abstract
With the rapid advancement of Large Models, numerous text-and-vision-fused Multimodal Large Models (MLMs) have emerged. However, these MLMs remain susceptible to informational interference in visual perception, particularly in color perception, which introduces an additional risk of hallucination. To validate this hypothesis, we introduce the ”What Color Is It” dataset, a novel benchmark constructed using a simple method to trigger single-modality visual hallucination in MLMs. Based on this dataset, we further investigate the underlying causes of hallucination in the visual modality of MLMs and propose potential solutions to enhance their robustness.
I Introduction
Multimodal Large Language Models (MLLMs) are becoming increasingly sophisticated, driving continuous breakthroughs in fields such as image recognition. Their inherent architectural design provides robust advantages, particularly in terms of generalizability and zero-shot performance across diverse tasks. However, a contradiction has emerged in the literature: despite their prowess in high-level visual understanding, the most current MLLMs exhibit persistent and notable shortcomings in the elemental functions of color identification and perception. This finding underscores the need to address fundamental perceptual limitations before MLLMs can achieve true human-like visual intelligence.
The rigorous evaluation of color perception in MLLMs necessitates specialized, high-quality test data. Addressing this need, the recently proposed ColorBench dataset offers a rich corpus designed for an all-encompassing measurement of models’ color sensibility. While instrumental, this benchmark overlooks a crucial aspect of multimodal processing. MLLMs are susceptible to semantic leakage where surrounding contextual information, such as descriptive text or explicit color terminology present in the images, can interfere with their visual judgment. Failure to isolate and mitigate this semantic interference risks producing inflated performance metrics that do not truly reflect the models’ underlying capacity for basic color perception.
Figure 1: Basic Structure of Samples in the ”What color is it” Dataset. The samples consist of three core components (background, question, and text), where text is the key element. Multimodal models need to identify the visual color of the text based on the guidance of the question in the image. The text can be color-irrelevant words or color-related words (serving as distractors), while the background and question can be black-and-white or colored (as distractors) as needed.
In the context of multimodal instructional hallucination, pioneering research [zhang2025modalities] has investigated MLLMs’ response bias and certainty when the input text and image modalities present a direct conflict. Our work significantly advances this line of inquiry by modifying the locus of the conflict. Instead of relying on a separate conflicting text prompt, we embed the textual instruction and the conflicting cue directly within the image plane, as shown in Figure 1. The stimuli we design are composed of three integrated components: the background visual cue, the query text, and the interfering textual color name. This fusion ensures that the interference (e.g., the visual color of the background, the text color of the question, and the explicit color name text) and the core visual data reside within the same modality. The model is thus forced to resolve the discrepancy between these visual-semantic elements, allowing us to thoroughly explore the resulting hallucination phenomena. This distinct approach contrasts sharply with previous studies, as we intentionally shift the creation of conflicting information from the external textual modality to the internal, complex image modality.
Our main contributions are summarized as follows.
  • Pioneering Dataset Construction: While the data construction methodology, such as ”What color is it,” appears straightforward and a few similar samples might be incidentally included in existing datasets, we are the first to systematically generate and thoroughly investigate such data, resulting in the creation of a novel dataset.
  • Exploring Multi-Modal Hallucination: We extend this dataset by varying the color description in the prompt and introducing a text masking mechanism. This allows us to systematically explore the influence of color-query question on the hallucination phenomenon in MLLMs.
  • Extensive Model Evaluation and Analysis: We rigorously test a wide array of state-of-the-art multimodal models on our proposed dataset. Furthermore, we conduct detailed qualitative and quantitative case analysis on the results from reasoning-capable MLLMs, from which we derive key findings and conclusions.
II Related Work
II-A Multimodal Large Language Models Hallucination
Hallucination in Multimodal Large Language Models (MLLMs) is a critical robustness issue, defined as generating undesirable outputs that are fundamentally unfaithful to the visual input, even if linguistically fluent. The most prevalent form is Object Hallucination (OH), characterized by predicting non-existent or inaccurate objects and categorized by category, attribute, or relation errors[dai2023plausible]. This phenomenon often stems from an over-reliance on the powerful language model’s linguistic prior, exacerbated by the imbalance between pure textual and multi-modal training data, a tendency sometimes termed ”blind faith in text”. Evaluation relies on visual fidelity metrics like CHAIR, which quantifies the proportion of ungrounded object words. Mitigation strategies focus on enforcing tighter visual grounding, such as the proposed Object-Masked Language Modeling loss, which enhances token-level image-text alignment to significantly reduce OH rates.
Hallucination in Multimodal Large Language Models (MLLMs) is a critical robustness issue, defined as generating undesirable outputs that are fundamentally unfaithful to the visual input, even if linguistically fluent[bai2024hallucination, sun2024aligning]. The most prevalent form is Object Hallucination (OH), characterized by predicting non-existent or inaccurate objects and categorized by category, attribute, or relation errors[dai2023plausible, chen2024multi]. This phenomenon often stems from an over-reliance on the powerful language model’s linguistic prior, exacerbated by the imbalance between pure textual and multi-modal training data, a tendency sometimes termed ”blind faith in text”[rohrbach2018object]. Several methods have been suggested to mitigate the object hallucination issue, such as integrating an external object detector[shang2025pixels], applying visually grounded visual instruction tuning[zhang2024groundhog] or reinforcement learning[gunjal2024detecting],performing iterative refinement[zhou2023analyzing], and adapting the decoding strategies[huang2024opera].
II-B Color Detection and Benchmarking in MLLM
MLLM evaluation has shifted from generalized accuracy towards diagnostic and specialized benchmarks that probe fine-grained perception and robustness failures in complex scenarios[ye2025assessing]. Specialized platforms like VisualAgentBench (VAB) assess models as interactive visual agents across domains such as Embodied and Graphical User Interface (GUI) tasks[gomez2025color], while OCRBench rigorously evaluates capabilities in comprehending text-rich visual data, including multilingual and mathematical expressions. Furthermore, dedicated benchmarks are crucial for assessing low-level visual fidelity, exemplified by COLORBENCH[liang2025colorbench], which provides a granular view of model performance across 11 tasks spanning Color Perception, Reasoning (e.g., Color Extraction, Color Proportion), and Robustness. Adversarial robustness is specifically tested by ColorBlindnessEval[ling2025colorblindnesseval], which uses Ishihara-like patterns to challenge numerical recognition under complex visual contexts, revealing significant performance drops due to susceptibility to visual noise.
II-C Cross-modal Information Conflict in MLLM
Cross-modal information conflict occurs when Multimodal Large Language Models (MLLMs) receive contradictory facts across modalities[yan2025multimodal], revealing a susceptibility often manifested as ”blind faith in text”—a systemic tendency to favor text over visual evidence[zhu2024unraveling]. Recent research formalizes this decision process, termed modality following, by decomposing it into two governing principles: Relative Reasoning Uncertainty ($\Delta H_{rel}$), measured by the confidence gap (entropy) between unimodal inputs, and a stable Inherent Modality Preference, quantified by the unique Balance Point where uncertainties are perceived as equal[hayashi2025diagnosing]. This framework revealed that in ambiguous regions, models exhibit internal layer-wise oscillations between conflicting answers[elazar2022measuring]. Mitigation techniques include inference-time interventions like Dynamic Contrastive Decoding (DCD), which removes undesirable logits from the less confident modality, and approaches utilizing Representation Engineering to probe and steer modality preference directly in the latent space[wang2024mdpo].
Moreover, studies show that models often favor one modality over another, such as prioritizing the image despite contradictory captions, with different models varying in which modality they favor[hua2025vision]. This preference is reflected in the model’s internal structure, where specific attention heads can be adjusted to favor one modality. Additionally, modality-agnostic ”router heads” can be leveraged to improve performance across different modalities and datasets.
III Benchmark Design
The dataset designed in this study consists of three main subsets, namely the color subset, simple subset, and mask subset. Each subset is further divided into four types of samples, whose specific characteristics are detailed as follows:
  1. 1.
    Type 1: Only the text exhibits specific color characteristics;
  2. 2.
    Type 2: In addition to the text, the question element also has color characteristics, which serve as interference;
  3. 3.
    Type 3: On the basis of Type 1, the recognition difficulty is further increased—each character of the text is assigned a unique color;
  4. 4.
    Type 4: Based on Type 2, an additional type of text is introduced and endowed with another color that is different from the color of the previous text.
For each of the aforementioned sample types, 100 images have been generated for model testing.
The core function of the color subset is to interfere with the multimodal large model’s judgment on the true color of the current text by setting the content of the text component in the image to a color name different from its actual color, as specifically illustrated in Figure 2. Its generation logic is as follows: For Type 1(Figure 2(a)), a specific color name is first selected from a color library (shown in Appendix A) as a distractor for the text component, and then a color irrelevant to the distractor is chosen from the same color library as the actual color of the text (i.e., the correct answer)—here, ”irrelevant to the distractor” does not merely mean the two colors are different, but rather requires that their similarity ($\Delta E$) is lower than a preset threshold; Type 2(Figure 2(b)) further selects a color irrelevant to both the distractor and the correct answer as the color specified in the question on the basis of Type 1; Type 3(Figure 2(c)) screens out n colors irrelevant to the distractor from the color library to form the actual color set of the text (i.e., the correct answer set); for Type 4(Figure 2(d)), two specific color names are first selected from the color library (shown in Appendix A) as distractors for the text component, followed by selecting two colors irrelevant to the distractors from the color library as the actual colors of the text respectively, and finally an additional color irrelevant to both the distractors and the correct answers is chosen as the color specified in the question.
(a) type-1
(b) type-2
(c) type-3
(d) type-4
Figure 2: Sample Image Examples of the ”Color” Subset. Type 1 modifies only the color of the text, with the background and question remaining black-and-white; Type 2 alters the colors of both the text and the question, while the background stays white; Type 3 changes the color of each character in the text, with the background and question kept black-and-white; Type 4 adjusts the colors of two words in the text and the question, and the background remains white.
The Simple subset serves as the control group, adopting several color-irrelevant words (e.g., Saturday, good; specific examples are provided in Appendix B) as text to avoid the interference of text semantic information on text color recognition in images, with the specific setup illustrated in Figure 3. This dataset is designed to verify whether multimodal models possess the ability to recognize text colors in images.
(a) type-1
(b) type-2
(c) type-3
(d) type-4
Figure 3: Sample Image Examples of the ”Simple” Subset. The color processing method for each type is consistent with that of the ”Color” Subset, except that the words in the text are replaced from color-related ones to those completely irrelevant to color.
The generation method of the mask subset is roughly consistent with that of the color subset. The only difference lies in the process of image rendering: by randomly masking some characters in the text (the number of masked characters does not exceed half of the text length), the interference caused by the color text attribute of the text is weakened, as specifically shown in Figure 4. The setting of the mask dataset is mainly to explore whether the multimodal model will still generate responses interfered by the semantic information of the text after the interference of the semantic modality is weakened.
(a) type-1
(b) type-2
(c) type-3
(d) type-4
Figure 4: Sample Image Examples of the ”mask” Subset. The color processing method for each type is consistent with that of the ”Color” Subset, except that partial characters in the text are randomly masked with ”*”.
IV Experiments and Results
IV-A Experimental Setup
To comprehensively evaluate the dataset we constructed, we conducted tests on three subsets using 12 models, including 6 accessed via API calls and 6 deployed locally such as Grok, GPT, Gemini, Qwen and others. Our test scope also includes the latest reasoning models, whose output results will be discussed in detail in the analysis section. Regarding prompt design, since all distractors are within the visual modality, we did not add additional distractions or prompts to the text-based questions and only instructed the multimodal models to answer the questions presented in the images.
IV-B Metrics
Given that the dataset we designed covers 4 categories of data, and the interference items include the textual information in the text and the color information of the Question, we have formulated the following indicators to measure the multimodal hallucination performance of multimodal models on such tasks.
IV-B1 Recognition Accuracy
A model’s answer is deemed correct if it can accurately output the true color of the text instead of the color mentioned in the textual information. This judgment criterion is consistent with the response logic of human testers when solving such image-based questions. The formula for Recognition Accuracy is defined as follows.
Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question, and $A$ be the answer provided by $M$. We define $A$ as ”correct” if and only if:
$A\text{ is correct}\iff A=\text{VisualColor}(\text{Text}(I))$
where $\text{Text}(I)$ extracts the text content from $I$, and $\text{VisualColor}(T)$ denotes the visual color of the text $T$.
IV-B2 Hallucination 1 — Hallucination Caused by Text Information Interference
A model is considered to have triggered the first type of hallucination (i.e., hallucination caused by text information interference) if it fails to output the true color of the text and directly takes the color contained in the textual information as the final result. The formula for this type of hallucination is defined as follows.
Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question about text color, and $A$ be the answer (color) provided by $M$. The function $\text{Text}(I)$ extracts the text string in $I$. The function $\text{VisualColor}(\text{Text}(I))$ is the true visual color of the text in $I$. The set $\mathcal{S}(\text{Text}(I))$ is the **semantic color set** (i.e., colors commonly associated with the text content).
The model $M$ exhibits a ”Type I Hallucination” if and only if its answer $A$ satisfies the following two conditions:
$\text{Type I Hallucination}\iff A\in\mathcal{S}(\text{Text}(I))$
IV-B3 Hallucination 2 — Hallucination Caused by Question Color Interference
A model is considered to have triggered the second type of hallucination (i.e., hallucination caused by question color interference) if it fails to output the true color of the text and selects the color of the Question text as the final result. The formula for this type of hallucination is defined as follows.
Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question, and $A$ be the color answer provided by $M$. The function $\text{Text}(I)$ extracts the text string in $I$. The set $\mathcal{S}(Q)$ is the **semantic color set** indicated by the content of the question $Q$.
The model $M$ exhibits a ”Type II Hallucination” if and only if its answer $A$ satisfies:
$\text{Type II Hallucination}\iff A\in\mathcal{S}(Q)$
Due to slight differences in the design purposes and generation methods of the three subsets, the indicators adopted in the experimental tests also vary. The specific indicator configuration methods are shown in Table I. It is worth noting that the dataset has avoided the repetition between interference items and the true color of the text during generation. Therefore, any color output by the multimodal model that is included in the interference items is regarded as triggering hallucination.
TABLE I: Metrics Selection for the Subsets of ”What color is it”.
dataset task Acc Hallu-1 HAllu-2
Color type-1
type-2
type-3
type-4
Simple type-1
type-2
type-3
type-4
Mask type-1
type-2
type-3
type-4
IV-C Results
The model test results on the Color, Simple, and Mask subsets are summarized in Table II, Table III, and Table IV, respectively. Notably, since the Simple subset contains no interference information, the phenomenon of Hallucination 1 does not occur. Therefore, when calculating the performance metrics, we only consider the Accuracy and the Hallucination Rate for Hallucination 2 (specifically the type caused by problematic color interference).
TABLE II: Evaluation Results of Multimodal Models on the color Sub-dataset.
Type-1 Type-2 Type-3 Type-4
Model acc h1 acc h1 h2 acc h1 acc h1 h2
grok-4 43 25 50 21 2 0 70 11 70 0
gpt-5 42 29 62 12 3 2 13 23 42 2
gpt-4o 55 15 62 4 8 6 5 9 20 2
gemini2.0-flash 31 56 7 24 58 3 28 5 45 39
qwen3-vl-plus-thinking 24 59 44 35 2 1 9 10 75 0
qwen3-vl-plus 5 93 0 97 2 1 12 0 98 0
qwen3-vl-32b 1 97 0 93 4 1 26 0 100 0
internvl3.5-14B 0 100 0 100 0 0 64 0 100 0
glm4v-9B 0 76 0 76 0 0 21 0 48 0
qwen3-vl-8B 0 99 0 99 1 0 34 0 100 0
qianfan-vl-8B 0 100 0 100 0 0 92 0 100 0
qwen2.5-vl-7B 0 100 0 100 0 0 73 0 100 0
TABLE III: Evaluation Results of Multimodal Models on the simple Sub-dataset.
Type-1 Type-2 Type-3 Type-4
Model acc acc h2 acc acc h2
grok-4 52 46 5 1 23 8
gpt-5 58 61 2 4 33 4
gpt-4o 64 54 13 5 3 1
gemini2.0-flash 75 10 74 6 3 72
qwen3-vl-plus-thinking 50 49 1 6 13 7
qwen3-vl-plus 54 29 12 7 1 15
qwen3-vl-32b 50 16 46 4 9 42
internvl3.5-14B 39 27 8 0 4 34
glm4v-9B 36 20 3 0 0 1
qwen3-vl-8B 53 26 24 2 2 33
qianfan-vl-8B 40 23 6 1 16 10
qwen2.5-vl-7B 31 9 34 3 8 32
TABLE IV: Evaluation Results of Multimodal Models on the mask Sub-dataset.
Type-1 Type-2 Type-3 Type-4
Model acc h1 acc h1 h2 acc h1 acc h1 h2
grok-4 46 31 31 39 1 1 21 7 79 0
gpt-5 24 67 23 71 1 1 18 1 95 0
gpt-4o 36 47 36 36 10 4 9 12 66 0
gemini2.0-flash 60 17 22 3 58 0 8 9 18 49
qwen3-vl-plus-thinking 31 53 26 57 2 1 16 12 59 1
qwen3-vl-plus 44 35 27 40 9 1 13 6 46 1
qwen3-vl-32b 33 44 24 39 15 1 21 5 73 4
internvl3.5-14B 36 33 28 36 8 0 27 2 57 11
glm4v-9B 24 18 13 18 1 0 2 0 23 0
qwen3-vl-8B 51 23 30 24 17 0 17 4 53 7
qianfan-vl-8B 30 48 20 33 5 0 38 9 60 7
qwen2.5-vl-7B 16 57 10 49 10 0 20 5 63 6
V Analysis
We performed both statistical analysis and case studies on the obtained experimental results. The statistical analysis, primarily based on the recognition accuracy and the trigger rates of Hallucination 1 and Hallucination 2 for various models presented in Table II, Table III, and Table IV, provides a preliminary assessment of the overall performance of the multimodal models on this dataset. The case studies, focusing on typical correct and incorrect samples, offer an in-depth analysis of the Chain-of-Thought (CoT) reasoning process employed by the two large multimodal models incorporating advanced inference mechanisms.
V-A Statics Analysis
We first analyze the results presented in Table II, where we found that most Vision-Language Models (VLMs) with a small number of parameters exhibited severe hallucinations (Hallu-1 and Hallu-2) during testing on the Color subset. Although these two types of hallucinations are alleviated with the increase in model parameters and the introduction of advanced reasoning mechanisms, they still persist.
Next, a comparative analysis of Table II and Table IV reveals that randomly masking certain characters in the interfering text weakens the color semantic information contained within it. Consequently, the hallucination rate of most VLMs on the Mask subset is reduced, although Hallu-1 type hallucinations are still observed.
Specifically, the Hallucination 1 rate for Grok-4,GPT-5 and GPT-4o on the Mask subset unexpectedly increased. We hypothesize two potential reasons:
  1. 1.
    Their pre-training data might contain samples similar to those in the Color subset, whereas the Mask subset data, due to the presence of masks, may have been excluded from the training corpus.
  2. 2.
    They might prioritize treating this type of question as a word-filling (or in-filling) task and directly outputting the filled-in result just like V-B2. Since this behavior deviates from the task objective, it is still categorized as a Type 1 hallucination.
Finally, comparing the experimental results inTable II and Table III, we observe a further improvement in the VLM’s accuracy on the Simple subset when the interfering text is completely replaced with non-interfering text (i.e., text entirely devoid of color semantic information). This demonstrates the VLM’s capability to comprehend instructions within the image and provide the correct answer (i.e., the text’s actual color). However, color recognition errors still occur on some samples, which aligns with the phenomena reported in the study by [liang2025colorbench]. This suggests that the current capability of VLMs in the domain of color perception remains relatively weak and requires further improvement.
V-B Cases Analysis
This study evaluates the performance of two prominent multi-modal large language models (MLLMs) equipped with reasoning capabilities: GLM4v-Thinking and Qwen3-VL-Thinking. We specifically analyze a set of representative reasoning cases from the Color and Mask subsets, aiming to delve into the underlying causes of the models’ reasoning successes and failures.
V-B1 GLM4v-Thinking
First, by comparing the test results of the GLM4v-Thinking model on the Color and Simple subsets (as shown in Figure 5(a) and Figure 5(b), respectively), we observed that, despite the MLLM correctly identifying the color of the Text in the initial reasoning stage, the interfering textual information in the Color subset still led the GLM4v-Thinking model to select the color represented by the Text (the distractor color) as the final reasoning result in the subsequent stages. Conversely, on the Simple subset, where no interfering text was present, the GLM4v-Thinking model was able to successfully complete the entire reasoning process. This phenomenon suggests that intra-modal conflicting information can also influence the reasoning process of MLLMs, consequently leading to hallucinations.
(a) Test Case on the Color Subset
(b) Test Case on the Simple Subset
Figure 5: Test Cases of the GLMv4-Thinking Model on the Color and Simple Subsets. Red parts indicate hallucinatory reasoning content, while green parts indicate correct reasoning content.
Further testing on the Masked subset revealed that the GLM4v-Thinking model adopts distinct processing strategies for different types of masked data. Specifically, as shown in Figure 6(a), the model automatically completes the masked information within the image and erroneously utilizes the interfering color information introduced by this completion as the inference result. Conversely, in Figure 6(b), the model elects to treat the masked text as a random string and successfully extracts the correct color information as the final inference result. This observation clearly indicates that the inference model still carries a risk of hallucination when handling these types of tasks.
(a) Incorrect Test Case on the Mask Subset
(b) Correct Test Case on the Mask Subset
Figure 6: Test Cases of the GLMv4-Thinking Model on the mask Subset. Red parts indicate hallucinatory reasoning content, while green parts indicate correct reasoning content.
Finally, a case study analyzing the Type 3 and Type 4 samples of the Color subset for the GLM4v-Thinking model (as illustrated in Figure 7 and Figure 8) revealed the following: In complex scenes involving multiple mixed colors, the inference model is initially capable of correctly identifying almost every color. However, the increased color complexity does not mitigate the interference from the text information itself, and the model ultimately selects the interfering color information as the final inference result in the later stages of reasoning.
Figure 7: Test Cases of the GLMv4-Thinking Model on the Color Subset (Type-3). Red parts indicate hallucinatory reasoning content, while green parts indicate correct reasoning content.
Figure 8: Test Cases of the GLMv4-Thinking Model on the Color Subset (Type-4). Red parts indicate hallucinatory reasoning content, while green parts indicate correct reasoning content.
V-B2 Qwen3V-plus-thinking
Similarly, through an in-depth analysis of the test cases for Qwen3-VL-Thinking on the Color (Figure 9) and Mask (Fig. 10) subsets, we observe that the cross-modal oscillation phenomenon, as documented in XXX’s paper, can also occur within a single modality when conflicting information is present.
(a) Incorrect Test Case on the Color Subset
(b) Correct Test Case on the Color Subset
Figure 9: Test Cases of the Qwen3-VL-Plus-Thinking Model on the Color Subset. Red parts indicate hallucinatory reasoning content, while green parts indicate correct reasoning content.
Specifically, the reasoning process on the Color subset shows noticeable fluctuations. In the case of Figure 9(a), the reasoning model tends to prioritize the distracting color information; conversely, in Figure 9(b), it manages to focus on the true color information in the text and successfully identifies the potential conflicting information trap.
(a) Incorrect Test Case on the Mask Subset
(b) Correct Test Case on the Mask Subset
Figure 10: Test Cases of the Qwen3-VL-Plus-Thinking Model on the Mask Subset. Red parts indicate hallucinatory reasoning content, while green parts indicate correct reasoning content.
For the Mask subset test cases, although Qwen3-VL-Thinking initially treats the Text with Mask as an information completion problem in the early stage of reasoning, subsequent reasoning exhibits two distinct tendencies: either prioritizing the distracting color information (Figure 10(a)) or focusing on the true color information of the Text (Figure 10(b)), resulting in completely different outcomes.
Furthermore, the recurrent ”wait” oscillation behavior observed in the Chain-of-Thought for the Type 5 task case of the Color subset (illustrated in Figure 12) strongly corroborates this finding.
Figure 11: Test Cases of the Qwen3-VL-Plus-Thinking Model on the Color Subset (Type-3). Red parts indicate hallucinatory reasoning content, while green parts indicate correct reasoning content.
Finally, in the test case for the type4 task of the Color subset (as shown in Figure 11), the Qwen3-VL-Thinking model, when identifying the color of the second character, was directly influenced by the interfering color information and erroneously designated it as the character’s color, despite a significant color discrepancy between the two. This finding suggests that an information conflict arising within a single image modality can substantially impact the results of even fundamental tasks, such as image color recognition.
Figure 12: Test Cases of the Qwen3-VL-Plus-Thinking Model on the Color Subset (Type-4). Red parts indicate hallucinatory reasoning content, while green parts indicate correct reasoning content.
VI Conclusion
We designed the ”What color is it” dataset to specifically tackle the observed deficiencies in color perception within current multimodal large language models (MLLMs). Through a series of rigorous experiments and comprehensive testing, we confirmed that, even when constrained to the unimodal domain of image processing, the presence of contradictory visual cues causes contemporary MLLMs to prioritize textual information embedded within the image over the core visual information. This hierarchical or priority-based information processing can subsequently induce ”loyalty hallucinations” in the MLLMs.

### v2

@@
 The formula for Recognition Accuracy is defined as follows.
-Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question, and $A$ be the answer provided by $M$. We define $A$ as ”correct” if and only if:
+Let $M$ be a multimodal model, II be an input image, $Q$ be a question, and AA be the answer provided by $M$. We define AA as ”correct” if and only if:
 $A\text{ is correct}\iff A=\text{VisualColor}(\text{Text}(I))$
-where $\text{Text}(I)$ extracts the text content from $I$, and $\text{VisualColor}(T)$ denotes the visual color of the text $T$.
+where $\text{Text}(I)$ extracts the text content from II, and $\text{VisualColor}(T)$ denotes the visual color of the text $T$.
 IV-B2 Hallucination 1 — Hallucination Caused by Text Information Interference
@@
 The formula for this type of hallucination is defined as follows.
-Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question about text color, and $A$ be the answer (color) provided by $M$.
-The function $\text{Text}(I)$ extracts the text string in $I$.
-The function $\text{VisualColor}(\text{Text}(I))$ is the true visual color of the text in $I$.
+Let $M$ be a multimodal model, II be an input image, $Q$ be a question about text color, and AA be the answer (color) provided by $M$.
+The function $\text{Text}(I)$ extracts the text string in II.
+The function $\text{VisualColor}(\text{Text}(I))$ is the true visual color of the text in II.
 The set $\mathcal{S}(\text{Text}(I))$ is the **semantic color set** (i.e., colors commonly associated with the text content).
-The model $M$ exhibits a ”Type I Hallucination” if and only if its answer $A$ satisfies the following two conditions:
+The model $M$ exhibits a ”Type I Hallucination” if and only if its answer AA satisfies the following two conditions:
 $\text{Type I Hallucination}\iff A\in\mathcal{S}(\text{Text}(I))$
@@
 The formula for this type of hallucination is defined as follows.
-Let $M$ be a multimodal model, $I$ be an input image, $Q$ be a question, and $A$ be the color answer provided by $M$.
-The function $\text{Text}(I)$ extracts the text string in $I$.
+Let $M$ be a multimodal model, II be an input image, $Q$ be a question, and AA be the color answer provided by $M$.
+The function $\text{Text}(I)$ extracts the text string in II.
 The set $\mathcal{S}(Q)$ is the **semantic color set** indicated by the content of the question $Q$.
-The model $M$ exhibits a ”Type II Hallucination” if and only if its answer $A$ satisfies:
+The model $M$ exhibits a ”Type II Hallucination” if and only if its answer AA satisfies:
 $\text{Type II Hallucination}\iff A\in\mathcal{S}(Q)$